Tyrell Co. entered into the following transactions involving short-term liabilities in 2017 and 2018.

2017

Apr. 20 Purchased $38,000 of merchandise on credit from Locust, terms n/30. Tyrell uses the perpetual inventory system.
May 19 Replaced the April 20 account payable to Locust with a 90-day, $35,000 note bearing 9% annual interest along with paying $3,000 in cash.
July 8 Borrowed $54,000 cash from NBR Bank by signing a 120-day, 10% interest-bearing note with a face value of $54,000.
__?__ Paid the amount due on the note to Locust at the maturity date.
__?__ Paid the amount due on the note to NBR Bank at the maturity date.
Nov. 28 Borrowed $30,000 cash from Fargo Bank by signing a 60-day, 7% interest-bearing note with a face value of $30,000.
Dec. 31 Recorded an adjusting entry for accrued interest on the note to Fargo Bank.

2018

__?__ Paid the amount due on the note to Fargo Bank at the maturity date.

Required information [The following information applies to the questions displayed below.) Tyrell Co. entered into the followRequired information (The following information applies to the questions displayed below.] Tyrell Co. entered into the followRequired information [The following information applies to the questions displayed below.] Tyrell Co. entered into the follow5.1 Prepare journal entries for all the preceding transactions and events for 2017. (Do not round your intermediate calculatiView transaction list Journal entry worksheet 1 3 4 5 6 7 Replaced the April 20 account payable to Locust with a 90-day, $35,5.1 Prepare journal entries for all the preceding transactions and events for 2017. (Do not round your intermediate calculatiJournal entry worksheet < 1 2 3 5 6 7 Paid the amount due on the note to Locust at the maturity date. Note: Enter debits befoJournal entry worksheet < 1 2 3 4 6 7 Paid the amount due on the note to NBR Bank at the maturity date. Note: Enter debits beJournal entry worksheet < 1 2 3 4 5 6 7 Borrowed $30,000 cash from Fargo Bank by signing a 60-day, 7% interest- bearing noteJournal entry worksheet < 1 2 3 4 5 6 7 Recorded an adjusting entry for accrued interest on the note to Fargo Bank. Note: EntJournal entry worksheet Paid the amount due on the note to Fargo Bank at the maturity date. Note: Enter debits before credits
Answer #1
1) Maturity date
locust NBR fargo
date of the note 19-May 8-Jul 28-Nov
term of note 90 120 60
maturity date 17-Aug 5-Nov 27-Jan
2) interest due at maturity
principal * Rate * time = interest
locust 35,000 * 9% * 90/360 = 788
NBR 54,000 * 10% * 120/360 = 1800
Fargo 30,000 * 7% * 60/360 = 350
3) Amount in adjusting entry
Fargo Bank
principal * Rate * time = interest
interest to be acccrued in 2016 30,000 * 7% * 33/360 = 193
4) interest expense to be recorded in 2017
principal * Rate * time = interest
interest to recorded in 2018 30,000 * 7% * 27/360 = 158
Journal entries
Date Accounting titles & Explanations Debit Credit
2016
20-Apr inventory 38,000
Accounts payable 38,000
19-May Accounts payable 38,000
cash 3,000
notes payable 35,000
8-Jul Cash 54,000
notes payable 54,000
17-Aug notes payable 35,000
interest expense 788
cash 35,788
5-Nov notes payable 54,000
interest expense 1,800
cash 55,800
28-Nov Cash 30,000
notes payable 30,000
31-Dec interest expense 193
interest payable 193
2017
27-Jan notes payable 30,000
interest payable 193
interest expense 158
cash 30,350